• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Chiranjib

    Log entry time 07:57:16 on August 21, 2006

    Entry number 181291

    keyword=OWL SHIFT SUMMARY

    Shift summary :::::::

    12 A ----- 1A :: Left arm Q1,Q3 was not working. Rob
    fixed that.

    1:20 A:::: asked for the beam 11 uA
    LD2 radiator 5%

    1:30A ::: 11 uA

    4:10 A ::: Radiator out LD2 18 uA

    4:40 A ::: Radiator in 5% LH2 11uA

    4:55 A ::: Asked for controlled access .

    Rob was trying to fix the magnet ( Q1)

    6 :00A ::: Asked for 11 uA on LH2 ( lets not call it
    for production)

    6:30 A :: Left arm Angle changed to 39.51.

    6:50 A : C optics radiator out

    7:00 A ::: 4 cm dummy radiator out

    7:10 A ::: 15 cm dummy radiator out

    7:25 A ::: tantalum radiator out

    7:35 A ::: 4 cm dummy radiator out



    *******************************************************************

    un Summary ::::::::


    RUN # Events (M)

    3016 1.9 on LD2 radiator 5%

    3017 1.9

    3018 1.9

    3019 1.9

    3020 1.9

    3021 1.9

    3022 1.9

    3023 1.9

    3024 1.9

    3025 1.9

    3026 1.9 on LD2 radiator out

    3027 637K on LH2 radiator 5%

    3028 99K on C optics radiator out

    3029 692K on 4 cm dummy radiator out

    3030 270K on 15cm dummy radiator out.

    3031 229K tentalum radiator out

    3032 on 4cm carbon